(4/y)-5=5/2y
Solve for y.

Found 2 solutions by tanimachatterjee, checkley71:
Answer by tanimachatterjee(60)

(4/y)-5=5/2y
Solve for y.
4/y - 5/2y = 5
or (8-5)/2y = 5 [adding two terms in rhs]
or 3 / 2y =5
or 3 = 10y
or 3/10 = y [ans]

Answer by checkley71(8403)

4/Y-5=(5/2)Y
4/Y-5Y/2=5
(8-5Y^2)/2Y=5
8-5Y^2=10Y
-5Y^2-10Y+8=0
5Y^2+10Y-8=0
USING THE QUADRATIC EQUATION WE GET
X=(-B+-SQRT[B^2-4AC])/2A
X=(-10+-SQRT[100-4*5*-8])/2*5
X=(-10+-SQRT[100+160)/10
X=(-10+-SQRT260)/10
X=(-10+-16.12)/10
x=(-10+16.12)/10
x=6.12/10
x=.612 solution.
x=(-10-16.12)/10
x=-26.12/10
x=-2.612 solution.